OpenStudy (morganchavez):

Happy Birthday! Or maybe Happy Un–Birthday? Either way, here is a fun birthday trick. Take the month you were born in. Multiply it by 4, add 13, multiply by 25, subtract 200, add the day you were born, multiply by 2, subtract 40, multiply by 50, add the last two digits of your birth year, subtract 10500…. viola! Your birth date! Impress your friends! Now, use your birthday for some more math fun. You can locate graph paper here. Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day.

OpenStudy (morganchavez):

Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Graph your function. Include your graph below. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. My birthday is July, 18, 1998.

OpenStudy (morganchavez):

@pkjag @One098 @monsterfox1997

OpenStudy (morganchavez):

@Niko_De_Angelo

OpenStudy (morganchavez):

ive worked on this all week n its over due bc i cnt do it ...n i cant afford an f in algebra 2 :"(

OpenStudy (one098):

@ganeshie8 Do you know this?

OpenStudy (one098):

Or know how to solve it?

OpenStudy (morganchavez):

f(x) = 1(x - 7)^2 + 18 to make it easy (the 1 is a, 7 is h and 18 is k) and then i would FOIL it out and get (x - 7)(x - 7) + 18 = x^2 - 14x + 49 + 18 = x^2 - 14x + 67 RIGHT?

OpenStudy (one098):

yes

OpenStudy (morganchavez):

and to find average rate of change from 4 to 7, plug in 7 for x then plug in 4 for x. Subtract the answers and put this over (7 - 4) = 3

OpenStudy (one098):

yes

OpenStudy (morganchavez):

but tht is the part where i am confused like i cant do the graph and tht

OpenStudy (morganchavez):

@just_one_last_goodbye @PaulaLovesSchool13

OpenStudy (anonymous):

Ur...what is the birthday part for?

OpenStudy (just_one_last_goodbye):

omg xD thought this was a birthday party in MAth section

OpenStudy (anonymous):

xD ^

OpenStudy (morganchavez):

i have no idea lol thts what i said .........and lol @just_one_last_goodbye

OpenStudy (morganchavez):

no but i seriously need help i did one part but the whol put 7 and x n all tht im lost

OpenStudy (just_one_last_goodbye):

im sorry @morganchavez im not good at math i know a few people that are tho ^_^ but im great at History and stuff so if u need help on that just tag me

OpenStudy (anonymous):

Same here, I can only do 6th grade or below math

OpenStudy (morganchavez):

ok can u tag those pple tht r please ive been on this for 3 days and it over due and dont i get it please

OpenStudy (morganchavez):

ok thank you @PaulaLovesSchool13

OpenStudy (anonymous):

Morgan, what REALLY was the birthday stuff for?

OpenStudy (morganchavez):

thats what the teacher gave it to us like that n we had to use the numbers from our bday meaning the month and day to put it in the equation ....ull learn this in 11 grade lol it is very difficult

OpenStudy (anonymous):

o_O teachers are becoming weirder and weirder

OpenStudy (morganchavez):

yasss and pain in tha butts

OpenStudy (anonymous):

And also, I am in 8th grade o_O

OpenStudy (morganchavez):

lol yup ur almost there

OpenStudy (just_one_last_goodbye):

@undeadknight26 @YanaSidlinskiy @texaschic101

OpenStudy (morganchavez):

thank u @just_one_last_goodbye ur awsome <3

OpenStudy (just_one_last_goodbye):

lol u are <3 ^_^

OpenStudy (yanasidlinskiy):

I'm NOT smart.

OpenStudy (just_one_last_goodbye):

-.= yesh u is

OpenStudy (anonymous):

you are a mathlete dude

OpenStudy (just_one_last_goodbye):

^ lol xD

OpenStudy (anonymous):

act like one o_O

OpenStudy (yanasidlinskiy):

fghabgruigba;kergbjaregba;gbiuargba;rgbrbg <------ That's what I got!!! I'm a Mathlete!!:)

OpenStudy (anonymous):

O_o

OpenStudy (anonymous):

Yana, come on

OpenStudy (anonymous):

unless u have to answer my math questions =_=

OpenStudy (yanasidlinskiy):

This is Random!! It's not really my Birthdate. My Birthdate is March. My birthday, just for practice: October 4, 1980 10 * 4 = 40 + 13 = 53 * 25 = 1325 – 200 = 1125 + 4 = 1129 * 2 = 2258 – 40 = 2218 * 50 = 110900 + 80 = 110980 – 10500 = 100480. The trick is that 100480 can be sub-divided down to 10 04 80, which is my birthday in the form (mm/dd/yy).

OpenStudy (anonymous):

o_O

OpenStudy (yanasidlinskiy):

OMg, lol!! Idk wut that is. It's only numbers....Lemme see if I can get a screenshot, from my word document. I had the same question before, Morgan. Found it in my Microsoft word!!:) Hang tight,xD!!

OpenStudy (anonymous):

average rate of change may mean the differential of your function \[\frac{df}{dx}\]

OpenStudy (morganchavez):

thank you sorry it took me forever to respond ....personL ISSUE feeling ill

OpenStudy (anonymous):

actually i'm the one who took forever, do you know how to differentiate a function?

OpenStudy (morganchavez):

no

OpenStudy (anonymous):

is this a homework/coursework question?

OpenStudy (morganchavez):

yes

OpenStudy (morganchavez):

@TheSmartOne @Teddyiswatshecallsme @almightyemery*** @boxhead06 @ameenbriscoe13 @ObeyBaby_nemo @J-bird @kavonta

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

do you how to graph a quadratic function using the values

OpenStudy (morganchavez):

im sure i culd figure it out probably

OpenStudy (morganchavez):

n hi

OpenStudy (anonymous):

hey morgan

OpenStudy (anonymous):

its your bday

OpenStudy (morganchavez):

no its a math assinment

OpenStudy (anonymous):

ooohhh

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

f(x) = 2(x - 6)^2 + 1 The vertex of this graph is at (6,1) If born 2 months earlier and 5 days later, (h,k) will be (4, 6) and f(x) will be: f(x) = 2(x - 4)^2 + 6 The vertex of this graph will be at (4,6) This will shift the graph 2 units to the left and 5 units up.

OpenStudy (anonymous):

just put in values of \[x=0\] to \[x\ge7\] in your function \[f(x)\]then the values you get for each \[x\] are a different \[y\]so you have to draw a table, then for all values \[(x,y)\]you get plot them on the graph

OpenStudy (morganchavez):

thank you @pkjag and @ameenbriscoe13

OpenStudy (anonymous):

wc

OpenStudy (anonymous):

anytime

OpenStudy (morganchavez):

@woooooooooowwwww

OpenStudy (anonymous):

medal please